OpenStudy (anonymous):

i need a refresher so that I can show my kids, can someone help me please.. 5x+y=18

OpenStudy (anonymous):

What do you need to do with it? Plot the line?

OpenStudy (anonymous):

find y

OpenStudy (anonymous):

Oh okay. Just subtract 5x from both side. Giving: \[5x+y-5x=18-5x \rightarrow y=-5x+18\]

OpenStudy (anonymous):

sides**

OpenStudy (anonymous):

thats it? That what I thought, but for some reason its looks wrong to me

OpenStudy (anonymous):

Thats all you have to do. You can do almost any manipulation as long as you do it to both sides. Anymore you want to look at? :P

OpenStudy (anonymous):

veron the method done by malebolence 19 is absolutely correct . i am also a kid of grade 8. tell ur kids to learn these type of methods . this would be helpful for them in future

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

No problem, any more troubles just post them :P

OpenStudy (anonymous):

ok rogster